OpenStudy (anonymous):

Evaluate the expression when a=3 and b=5 3/4a divided by 5b/6

OpenStudy (anonymous):

@Mr.ClayLordMath

OpenStudy (anonymous):

@SACAPUNTAS

OpenStudy (anonymous):

Evaluate the expression for a certain value means just put a 3 wherever you see \(a\) and put a 5 wherever you see \(b\). So if it's supposed to be \[\frac{3}{4a} / \frac{5b}{6}\] That just becomes \[\frac{3}{4(3)} / \frac{5(5)}{6}\]

OpenStudy (anonymous):

thx caN YOU HELP ME WITH ANOTHER QUESTION

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

A DEAD 53 1/2 foot tall tree is being cut down.on the first cut 6 2/8 are cut off.on the next cut,30 5/16 feet are cut off.how much of the tree remains to be cut off

OpenStudy (anonymous):

Unless I'm missing something that's just subtraction. \[53\frac{1}{2} - 6\frac{2}{8} - 30\frac{5}{16}\] Or if you prefer \[53\frac{1}{2} - (6\frac{2}{8} + 30\frac{5}{16})\]

OpenStudy (anonymous):

dont you have to simplifly

OpenStudy (anonymous):

Yes. I was just writing out the problem. You still have to actually do the subtraction.

OpenStudy (anonymous):

oh thx if theres any more problems i have trouble on i will @ you

OpenStudy (anonymous):

@SamSmiths